OpenStudy (anonymous):

Which value of x satisfies both 3x + y = 9 and 3x − 5y = 15? A) \[\frac{ 10 }{ 3 }\] B) \[\frac{ 2 }{ 3 }\] C) \[-1\] D) \[0\] @thomaster @satellite73 @maxdrizner

OpenStudy (anonymous):

3x + y = 9 --->(5) 3x - 5y = 15 ----------- 15x + 5y = 45 (result of multiplying by 5) 3x - 5y = 15 ------------add 18x = 60 x = 60/18 = 30/9 = 10/3 answer is 10/3

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

Nope, that definitely helped me A LOT! Thank you so much, you're an excellent help :).

OpenStudy (anonymous):

no problem :)